OpenStudy (anonymous):

Help!! For the angle 300 degrees find the six part: Sin, Cos, Tan, Csc, Sec, Cot in fraction form

OpenStudy (anonymous):

300 degrees is in the fourth quadrant. the equivalent angle in the first quadrant would be 360 - 300 = 60 degrees. this is a common angle that you can find the exact trigonometric functions for. they are: sin(60) = sqrt(3)/2 cos(60) = 1/2 tan(60) = sqrt(3) cot(60) = sqrt(3)/3 sec(60) = 2 csc(60) = 2*sqrt(3)/3